# Luminosity predictions for the first three ionisation stages of W, Pt and Au to probe potential sources of emission in kilonova

### v1

@@ 5 Conclusions
The suggested luminosities from observations of the AT2017gfo and the GRB230307A, from Hotokezaka et al. (2022) and Gillanders et al. (2023) respectively, have been used to determine mass estimates using the inverse procedure. Starting with the luminosity from observations, the required mass necessary to produce this value has been calculated using the available atomic data. For AT2017gfo the measured luminosity was erg s-1, which resulted in a mass of M⊙. For GRB230307A the measured luminosity was erg s-1, which resulted in a mass of M⊙. A total ejecta mass of M⊙ for AT2017gfo from Hotokezaka et al. (2022) implies that of the ejecta mass is W iii, while for GRB230307A a total ejecta mass of M⊙ from Gillanders et al. (2023) implies that of the ejecta mass is W iii. Assuming a similar amount of the first few ion stages this analysis would predict that of the AT2017gfo and of the GRB230307A are from W.
@@
In order to test plausibility the estimated amount of W iii is compared with the results from theoretical models of matter outflows of neutron-star mergers based on hydrodynamical simulations and nuclear-network calculations. These models, which resemble GW170817-like systems, yield total W masses in the range M⊙ to M⊙. Hence, broad consistency between the inferred masses from observations and theoretical predictions is found considering the involved uncertainties, which, on the theory side, include in particular the nuclear mass model employed in the nuclear network calculation. The mass determined from observation for AT2017gfo is close to the highest values predicted by our simulations, which is unsurprising due to the possibility that other lines or processes could be contributing to the feature in the observation. Concerning GRB230307A, the comparison reveals a somewhat worse match which may either be related to uncertainties of the inference or point to the ejection of more neutron-rich material in GRB230307A, for which no binary mass estimates exist since no gravitational waves were recorded from this event.
@@
Nuclear network calculations suggest that the production of W requires generally very neutron-rich conditions with an electron fraction , see Figure 8. Therefore, measurements of W are useful to estimate the amount of neutron-rich material from merger observations is therefore quite useful. Exploiting tight correlations between the W abundance and the mass fraction of lanthanides and actinides , we convert the inferred mass fraction of W to an estimate of . In the case of AT2017gfo, we find mass fractions of these groups of elements in the range 0.85 % to 11 % of the ejecta mass. These elements featuring a very high opacity are particularly relevant for shaping the light curve of KN and estimates from the light curve suggested tentatively lower values, while observations for metal-poor stars point to a high production in mergers if those are the dominant site of r-process nucleosynthesis. Similarly, W can be used to estimate the amount of third-peak r-process elements, which are constrained to the range 0.73% to 19% for AT2017gfo. It is therefore possible to infer from the observational estimate of the W mass in Section 3.2 that significant amounts of third-peak -process elements, lanthanides and actinides will be produced along with it. These considerations exemplify the usefulness of W to trace the properties of merger outflows and the underlying nucleosynthesis. It shows the importance of measuring W as laid out here and the need to improve observational data to more accurately determine the luminosity of features in KN spectra.

### v2

@@ 4 Implications of the limits on the ejected mass of W
5 Velocity distribution
A further test of consistency between merger models and observations is to consider the velocity distribution of the ejected in comparison to constraints on width of the observed spectral feature, which can be primarily attributed to Doppler broadening. Figure 11 shows the distribution of ejected mass versus radial velocity for four representative models from Table 4. In all these models, the ejection velocity spans a fairly wide range, peaking around . There is a noticeable difference, however, in the amount of low velocity () when comparing models with short-lived neutron-star remnants to those with long-lived remnants (the wind from a long-lived remnant leads to inner ejecta that are relatively -poor).
This difference in velocity distribution has implications for the expected profile shape which, given spectropscopic observations of sufficient quality, provides further constraints on the models. To illustrate this, we have computed simple optically-thin emission line profile shapes for the 4432.23 nm W iii transition (Figure 12) using the framework presented by Jerkstrand (2017), which is accurate to first order in . For these calculations we do not take into consideration any temperature/ionization variations in the ejecta, or any departures from spherical symmetry, and thus they should be considered only illustrative, pending more detailed modelling. The calculations assume that the emission is optically thin and that the emissivity is proportional to the product of the particle density and the total mass density (adopted as a proxy for the free-electron density, under the assumption of uniform ionization).
Despite the simplicity of the calculations, these profiles indicate the potential to obtain information on the distribution of W in the ejecta, and possibly the fate of the merger remnant, through kilonova observations. In particular, our models with short-lived neutron-star remnants suggest narrow profiles, owing to the relatively high densities at low velocity. In contrast, the absence of at low velocities leads to broad, somewhat flat topped profiles for our models with long-lived neutron-star remnants. More sophisticated modelling is required to draw firm conclusions from comparisons with the observations of AT2023vfi, but these simple calculations suggest somewhat better compatibility of the short-lived remnant models with the relatively narrow observed profile (FWHM Gillanders & Smartt 2024), although we note that the present calculation yields profiles that are somewhat too narrow. This may be indicative of line blending or a more extended velocity distribution, but more detailed modelling (involving temperature and ionization calculations) is needed to draw firm conclusions.
@@ 6 Conclusions
The suggested luminosities from observations of the AT2017gfo and the AT2023vfi, from Hotokezaka et al. (2022) and Gillanders & Smartt (2024) respectively, have been used to determine mass estimates using the inverse procedure. Starting with the luminosity from observations, the required mass necessary to produce this value has been calculated using the available atomic data. For AT2017gfo the measured luminosity was erg s-1, which resulted in a mass of M⊙. For AT2023vfi the measured luminosity was erg s-1, which resulted in a mass of M⊙. A total ejecta mass of M⊙ for AT2017gfo from Hotokezaka et al. (2022) implies that of the ejecta mass is W iii, while for AT2023vfi a total ejecta mass of M⊙ from Levan et al. (2024) implies that of the ejecta mass is W iii. Assuming a similar amount of the first few ion stages this analysis would predict that of the AT2017gfo and of the AT2023vfi are from W.
@@
To test plausibility the estimated amount of W iii is compared with the results from theoretical models of matter outflows of neutron-star mergers based on hydrodynamical simulations and nuclear-network calculations. These models, which resemble AT2017gfo-like systems, yield total W masses in the range M⊙ to M⊙. Hence, broad consistency between the inferred masses from observations and theoretical predictions is found considering the involved uncertainties, which, on the theory side, include in particular the nuclear mass model employed in the nuclear network calculation. The mass determined from observation for AT2017gfo is close to the highest values predicted by our simulations, which is unsurprising due to the possibility that other lines or processes could be contributing to the feature in the observation. Concerning AT2023vfi, the comparison reveals a somewhat worse match which may either be related to uncertainties of the inference or point to the ejection of more neutron-rich material in AT2023vfi, for which no binary mass estimates exist since no gravitational waves were recorded from this event.
@@
Nuclear network calculations suggest that the production of W requires generally very neutron-rich conditions with an electron fraction , see Figure 9. Therefore, measurements of W are useful to estimate the amount of neutron-rich material from merger observations is therefore quite useful. Exploiting tight correlations between the W abundance and the mass fraction of lanthanides and actinides , we convert the inferred mass fraction of W to an estimate of . In the case of AT2017gfo, we find mass fractions of these groups of elements in the range 0.85 % to 11 % of the ejecta mass. These elements featuring a very high opacity are particularly relevant for shaping the light curve of KN and estimates from the light curve suggested tentatively lower values, while observations for metal-poor stars point to a high production in mergers if those are the dominant site of r-process nucleosynthesis. Similarly, W can be used to estimate the amount of third-peak r-process elements, which are constrained to the range 0.73% to 19% for AT2017gfo. It is therefore possible to infer from the observational estimate of the W mass in Section 3.2 that significant amounts of third-peak -process elements, lanthanides and actinides will be produced along with it. Based on hydrodynamical merger models we estimate line profiles for the W feature, which are significantly affected by the tungsten production throughout the ejecta. This indicates the possibility to probe the velocity distribution of W and possibly distinguish short-lived and long-lived merger remnants if the latter indeed generically yield a strong deficiency of W at low velocities (v<0.1) as in our limited set of models. These considerations exemplify the usefulness of W to trace the properties of merger outflows and the underlying nucleosynthesis. It shows the importance of measuring W as laid out here and the need to improve observational data to more accurately determine the luminosity of features in KN spectra.
@@ Acknowledgements
Appendix A Effective Collision Strengths
The following tables contain effective collision strengths for the ions that were discussed in Section 3.1 and include the lines shown in Tables 1, 2 and 3. Tables 5, 6 and 7 contain W i, W ii and W iii. Tables 8, 9 and 10 contain Pt i, Pt ii and Pt iii. Tables 11, 12 and 13 contain Au i, Au ii and Au iii.
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 5986.99 | 1 | 2 | 5.32E-02 | 1.03E+00 | 1.02E+00 | 1.01E+00 | 1.01E+00 | 1.00E+00 |
| 6041.42 | 2 | 4 | 1.27E-01 | 2.25E+00 | 2.29E+00 | 2.33E+00 | 2.37E+00 | 2.39E+00 |
| 6646.86 | 4 | 5 | 1.35E-01 | 3.47E+00 | 3.47E+00 | 3.47E+00 | 3.47E+00 | 3.47E+00 |
| 7197.71 | 5 | 6 | 9.28E-02 | 3.64E+00 | 3.70E+00 | 3.75E+00 | 3.80E+00 | 3.83E+00 |
| 1272.63 | 2 | 7 | 8.07E+00 | 3.42E-01 | 3.29E-01 | 3.17E-01 | 3.06E-01 | 2.99E-01 |
| 3059.94 | 3 | 6 | 2.33E-03 | 3.25E+00 | 3.16E+00 | 3.08E+00 | 3.01E+00 | 2.96E+00 |
| 1682.76 | 6 | 8 | 1.33E+00 | 1.86E+00 | 1.68E+00 | 1.51E+00 | 1.36E+00 | 1.26E+00 |
| 1001.85 | 4 | 9 | 5.72E+00 | 3.15E-01 | 3.08E-01 | 3.02E-01 | 2.97E-01 | 2.93E-01 |
| 1173.91 | 5 | 10 | 1.66E+00 | 9.83E-01 | 9.02E-01 | 8.27E-01 | 7.63E-01 | 7.17E-01 |
| 1363.89 | 5 | 8 | 3.45E-01 | 8.80E-01 | 7.95E-01 | 7.17E-01 | 6.50E-01 | 6.02E-01 |
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 6584.02 | 1 | 2 | 2.06E-01 | 4.07E+00 | 3.76E+00 | 3.38E+00 | 2.99E+00 | 2.74E+00 |
| 6047.25 | 2 | 3 | 2.80E-01 | 5.41E+00 | 5.10E+00 | 4.69E+00 | 4.25E+00 | 3.94E+00 |
| 6477.50 | 3 | 4 | 1.86E-01 | 7.39E+00 | 6.91E+00 | 6.33E+00 | 5.77E+00 | 5.41E+00 |
| 6989.07 | 4 | 5 | 8.60E-02 | 7.99E+00 | 7.61E+00 | 7.11E+00 | 6.53E+00 | 6.10E+00 |
| 1367.26 | 2 | 8 | 2.81E+00 | 8.95E-01 | 8.50E-01 | 7.82E-01 | 6.99E-01 | 6.44E-01 |
| 1132.15 | 1 | 8 | 1.26E+00 | 7.20E-01 | 6.58E-01 | 5.71E-01 | 4.82E-01 | 4.26E-01 |
| 1347.66 | 1 | 6 | 2.21E-02 | 7.95E-01 | 7.72E-01 | 7.18E-01 | 6.49E-01 | 6.04E-01 |
| 1347.71 | 3 | 9 | 7.06E-01 | 1.96E+00 | 1.84E+00 | 1.65E+00 | 1.45E+00 | 1.31E+00 |
| 1694.50 | 2 | 6 | 1.93E-02 | 2.01E+00 | 1.95E+00 | 1.84E+00 | 1.69E+00 | 1.60E+00 |
| 1805.45 | 3 | 7 | 2.74E-02 | 1.88E+00 | 1.86E+00 | 1.75E+00 | 1.57E+00 | 1.44E+00 |
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 4432.23 | 1 | 2 | 5.04E-01 | 1.74E+00 | 1.58E+00 | 1.43E+00 | 1.30E+00 | 1.21E+00 |
| 4535.16 | 2 | 3 | 5.40E-01 | 4.59E+00 | 4.11E+00 | 3.67E+00 | 3.29E+00 | 3.02E+00 |
| 5504.75 | 3 | 4 | 2.60E-01 | 5.79E+00 | 5.28E+00 | 4.81E+00 | 4.41E+00 | 4.12E+00 |
| 1307.51 | 2 | 6 | 6.98E+00 | 6.23E-01 | 5.50E-01 | 4.83E-01 | 4.25E-01 | 3.84E-01 |
| 7097.87 | 4 | 5 | 7.13E-02 | 8.00E+00 | 7.21E+00 | 6.48E+00 | 5.84E+00 | 5.40E+00 |
| 1536.73 | 3 | 7 | 1.35E-01 | 1.81E+00 | 1.63E+00 | 1.47E+00 | 1.33E+00 | 1.23E+00 |
| 1187.67 | 3 | 9 | 3.60E+00 | 1.38E+00 | 1.24E+00 | 1.11E+00 | 9.98E-01 | 9.18E-01 |
| 804.69 | 1 | 8 | 9.31E-02 | 9.05E-01 | 8.10E-01 | 7.23E-01 | 6.47E-01 | 5.94E-01 |
| 1662.71 | 5 | 10 | 6.07E-01 | 4.79E+00 | 4.26E+00 | 3.77E+00 | 3.35E+00 | 3.05E+00 |
| 1147.80 | 2 | 7 | 3.00E-02 | 1.51E+00 | 1.36E+00 | 1.22E+00 | 1.10E+00 | 1.02E+00 |
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 1522.66 | 1 | 5 | 5.25E-01 | 2.69E+00 | 2.51E+00 | 2.34E+00 | 2.19E+00 | 2.09E+00 |
| 12888.66 | 1 | 2 | 4.09E-04 | 5.16E+00 | 4.88E+00 | 4.61E+00 | 4.38E+00 | 4.22E+00 |
| 1076.07 | 3 | 6 | 2.52E+00 | 9.62E-01 | 9.14E-01 | 8.70E-01 | 8.32E-01 | 8.05E-01 |
| 1726.65 | 2 | 5 | 1.40E-02 | 2.35E+00 | 2.19E+00 | 2.05E+00 | 1.93E+00 | 1.84E+00 |
| 1068.83 | 2 | 7 | 1.16E+00 | 8.73E-01 | 7.77E-01 | 6.89E-01 | 6.13E-01 | 5.59E-01 |
| 1864.18 | 2 | 4 | 4.36E-04 | 8.20E-01 | 7.40E-01 | 6.66E-01 | 6.02E-01 | 5.57E-01 |
| 740.95 | 1 | 8 | 9.59E-01 | 1.91E+00 | 1.80E+00 | 1.69E+00 | 1.59E+00 | 1.53E+00 |
| 1070.57 | 2 | 6 | 1.22E-01 | 1.30E+00 | 1.26E+00 | 1.22E+00 | 1.18E+00 | 1.16E+00 |
| 2805.51 | 5 | 7 | 1.32E-01 | 1.50E+00 | 1.41E+00 | 1.33E+00 | 1.26E+00 | 1.21E+00 |
| 1741.01 | 3 | 5 | 7.74E-05 | 2.02E+00 | 1.81E+00 | 1.61E+00 | 1.44E+00 | 1.32E+00 |
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 2188.35 | 2 | 4 | 2.54E+00 | 1.35E+01 | 1.24E+01 | 1.14E+01 | 1.05E+01 | 9.84E+00 |
| 1187.67 | 1 | 3 | 9.05E+00 | 1.82E+00 | 1.71E+00 | 1.61E+00 | 1.52E+00 | 1.46E+00 |
| 1068.80 | 1 | 4 | 1.10E-01 | 2.39E+00 | 2.28E+00 | 2.17E+00 | 2.07E+00 | 2.01E+00 |
| 2517.02 | 4 | 5 | 1.96E+00 | 4.49E+00 | 4.11E+00 | 3.76E+00 | 3.47E+00 | 3.25E+00 |
| 633.26 | 1 | 6 | 3.85E+00 | 1.41E+00 | 1.33E+00 | 1.25E+00 | 1.19E+00 | 1.14E+00 |
| 2036.89 | 3 | 5 | 2.68E-01 | 2.02E+00 | 1.86E+00 | 1.71E+00 | 1.58E+00 | 1.49E+00 |
| 2089.14 | 1 | 2 | 6.17E-06 | 3.07E+00 | 2.96E+00 | 2.86E+00 | 2.77E+00 | 2.71E+00 |
| 1339.65 | 4 | 7 | 6.46E+00 | 2.88E+00 | 2.65E+00 | 2.44E+00 | 2.26E+00 | 2.13E+00 |
| 751.25 | 2 | 8 | 1.52E+01 | 3.83E+00 | 3.46E+00 | 3.13E+00 | 2.85E+00 | 2.64E+00 |
| 750.23 | 1 | 5 | 3.93E-02 | 2.16E+00 | 2.05E+00 | 1.95E+00 | 1.86E+00 | 1.80E+00 |
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 1025.46 | 1 | 3 | 4.66E+01 | 4.27E+00 | 4.11E+00 | 3.97E+00 | 3.85E+00 | 3.76E+00 |
| 1126.28 | 2 | 4 | 2.22E+01 | 2.83E+00 | 2.60E+00 | 2.39E+00 | 2.21E+00 | 2.08E+00 |
| 2242.86 | 2 | 3 | 1.13E+00 | 3.82E+00 | 3.56E+00 | 3.33E+00 | 3.12E+00 | 2.98E+00 |
| 870.44 | 2 | 6 | 2.10E+01 | 1.61E+00 | 1.47E+00 | 1.34E+00 | 1.23E+00 | 1.15E+00 |
| 2262.34 | 3 | 4 | 3.05E+00 | 3.73E+00 | 3.43E+00 | 3.16E+00 | 2.93E+00 | 2.76E+00 |
| 468.81 | 1 | 7 | 1.91E+01 | 2.68E+00 | 2.65E+00 | 2.62E+00 | 2.60E+00 | 2.58E+00 |
| 3831.85 | 4 | 6 | 4.05E-01 | 2.12E+00 | 1.99E+00 | 1.88E+00 | 1.78E+00 | 1.71E+00 |
| 863.63 | 3 | 7 | 1.94E+00 | 4.02E+00 | 3.68E+00 | 3.38E+00 | 3.12E+00 | 2.93E+00 |
| 457.94 | 1 | 8 | 1.02E-02 | 3.11E+00 | 3.07E+00 | 3.04E+00 | 3.01E+00 | 2.99E+00 |
| 5427.12 | 5 | 6 | 2.00E-01 | 5.85E-01 | 5.37E-01 | 4.92E-01 | 4.53E-01 | 4.26E-01 |
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 1091.56 | 1 | 2 | 2.48E-02 | 1.55E+00 | 1.49E+00 | 1.45E+00 | 1.41E+00 | 1.38E+00 |
| 814.73 | 2 | 3 | 2.98E+01 | 6.45E-01 | 6.06E-01 | 5.70E-01 | 5.39E-01 | 5.17E-01 |
| 466.52 | 1 | 3 | 1.31E+00 | 1.10E+00 | 1.02E+00 | 9.48E-01 | 8.87E-01 | 8.44E-01 |
| 267.67 | 1 | 4 | 1.61E+08 | 8.30E-01 | 9.85E-01 | 1.13E+00 | 1.25E+00 | 1.34E+00 |
| 242.87 | 1 | 5 | 2.26E+08 | 9.96E-01 | 1.28E+00 | 1.55E+00 | 1.78E+00 | 1.94E+00 |
| 303.01 | 2 | 6 | 1.37E+05 | 2.34E+00 | 2.33E+00 | 2.31E+00 | 2.30E+00 | 2.29E+00 |
| 627.99 | 3 | 4 | 2.33E+06 | 4.18E-02 | 1.32E-01 | 2.16E-01 | 2.88E-01 | 3.39E-01 |
| 312.37 | 2 | 5 | 2.60E+07 | 2.60E-01 | 3.50E-01 | 4.34E-01 | 5.06E-01 | 5.57E-01 |
| 274.91 | 2 | 7 | 1.43E+06 | 4.47E+00 | 4.10E+00 | 3.76E+00 | 3.47E+00 | 3.26E+00 |
| 268.69 | 2 | 9 | 1.21E+06 | 2.10E+00 | 2.11E+00 | 2.13E+00 | 2.14E+00 | 2.15E+00 |
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 566.87 | 1 | 3 | 4.05E-01 | 5.29E-01 | 5.76E-01 | 6.18E-01 | 6.55E-01 | 6.81E-01 |
| 3844.62 | 2 | 3 | 2.87E-01 | 7.68E+00 | 7.61E+00 | 7.54E+00 | 7.48E+00 | 7.44E+00 |
| 987.64 | 3 | 4 | 2.72E+01 | 2.00E+00 | 1.78E+00 | 1.57E+00 | 1.40E+00 | 1.27E+00 |
| 685.79 | 2 | 5 | 2.71E+01 | 1.74E+00 | 1.60E+00 | 1.47E+00 | 1.36E+00 | 1.28E+00 |
| 337.60 | 1 | 5 | 8.27E+00 | 9.88E-01 | 9.16E-01 | 8.50E-01 | 7.92E-01 | 7.52E-01 |
| 834.68 | 3 | 5 | 1.74E+00 | 1.60E+00 | 1.50E+00 | 1.41E+00 | 1.34E+00 | 1.28E+00 |
| 785.78 | 2 | 4 | 4.98E-03 | 7.74E-01 | 6.96E-01 | 6.24E-01 | 5.62E-01 | 5.18E-01 |
| 5389.40 | 4 | 5 | 5.42E-02 | 5.07E+00 | 4.94E+00 | 4.83E+00 | 4.73E+00 | 4.66E+00 |
| 393.09 | 2 | 6 | 5.34E+00 | 2.09E+00 | 2.61E+00 | 3.09E+00 | 3.51E+00 | 3.80E+00 |
| 437.86 | 3 | 6 | 7.39E-01 | 1.61E+00 | 1.74E+00 | 1.85E+00 | 1.95E+00 | 2.02E+00 |
| Wavelength | Index | A-value | Effective Collision Strengths | |||||
| (nm) | Lower | Upper | (s-1) | 0.15 eV | 0.25 eV | 0.4 eV | 0.6 eV | 0.8 eV |
| 787.77 | 1 | 2 | 3.30E+01 | 3.39E+00 | 3.36E+00 | 3.33E+00 | 3.31E+00 | 3.29E+00 |
| 336.09 | 1 | 3 | 1.34E-01 | 2.66E+00 | 2.56E+00 | 2.47E+00 | 2.38E+00 | 2.33E+00 |
| 285.09 | 1 | 4 | 1.05E+01 | 2.39E+00 | 2.29E+00 | 2.20E+00 | 2.11E+00 | 2.06E+00 |
| 1878.69 | 3 | 4 | 3.72E+00 | 1.17E+01 | 1.10E+01 | 1.03E+01 | 9.73E+00 | 9.32E+00 |
| 257.58 | 1 | 5 | 7.95E+00 | 1.69E+00 | 1.59E+00 | 1.50E+00 | 1.43E+00 | 1.37E+00 |
| 247.86 | 1 | 6 | 2.04E+01 | 9.73E-01 | 9.49E-01 | 9.27E-01 | 9.07E-01 | 8.94E-01 |
| 382.73 | 2 | 5 | 1.09E+00 | 1.56E+00 | 1.37E+00 | 1.20E+00 | 1.05E+00 | 9.43E-01 |
| 446.77 | 2 | 4 | 2.78E-02 | 1.27E+00 | 1.17E+00 | 1.08E+00 | 1.01E+00 | 9.52E-01 |
| 2669.87 | 4 | 5 | 1.01E+00 | 5.10E+00 | 4.90E+00 | 4.73E+00 | 4.58E+00 | 4.47E+00 |
| 225.09 | 1 | 7 | 1.72E+01 | 1.11E+00 | 1.10E+00 | 1.08E+00 | 1.07E+00 | 1.06E+00 |